OpenStudy (anonymous):

A refrigerator manufacturer can sell all the refrigerators of a particular type that he can produce. The total cost of producing 'q' refrigerators per week is given by 300q+2000.the demand function is estimated as 500-2q. [a] Derive the revenue function. [b] Obtain the total profit function. [c] How many units per week should be produced in order to maximize profit? [d] What is the maximum profit available.

OpenStudy (dumbcow):

revenue = price*quantity p = 500-2q revenue-> R(q) = 500q -2q^2 profit = revenue -cost P(q) = R(q) -C(q) = (500q-2q^2)-(300q+2000) P(q) = -2q^2 +200q -2000

OpenStudy (anonymous):

thanks very much dumbcow.

OpenStudy (dumbcow):

yw about maximizing, find vertex of the parabola formed by P(q)

OpenStudy (anonymous):

ok.....

OpenStudy (dumbcow):

x = -b/2a max_q = -200/-4 = 50 plug that in to get max_profit

OpenStudy (anonymous):

am getting it

OpenStudy (anonymous):

pls dumbcow, how did u get the R[q]=500q-2q^2

OpenStudy (dumbcow):

oh sorry i multiplied p*q (500-2q)*q

OpenStudy (anonymous):

ok...thanks